how many milliliters of a 4% acid solution and how many milliliters of a 10% acid solution must be mixed to obtain 54mL of a 6% acid solution?
How do I solve?

Answer by richwmiller(17219)

x+y=54,
0.04*x+0.1*y=0.06*54
0.04*x+0.1*y=3.24
x=54-y
0.04*(54-y)+0.1*y=3.24
2.16-0.04y+0.1*y=3.24
0.06*y=1.08
y=18.0 at 10%
x=54-y
x=36.0 at 4%
check
0.04*36+0.1*18.0=0.06*54
1.44+1.8=3.24
3.24=3.24
ok